The area of a telescope lens is 7.195 x 10^3 mm^2

A) What is the area in square feet (ft^2)? Enter your answer in scientific notation

B) If it takes a technician 47.8 seconds to polish 1.57 x 10^2 mm^2, how long does it take her to polish the entire lens?
Answer #1

A) 1 mm​​​​2 = 1.076 × 10-5 ft​​2

Area in square feet =7.195 × 1.076 × 10-5 × 103 ft​​​​2

= 7.742 × 10-2 ft​​​​​2

B) 1.57 × 102 mm​​​​​2​ takes 47.8 seconds to polish.

7.195 × 103 mm​​​​​2​​​​ takes 47.8×7.195×103 /1.57×102 seconds

= 2190.6 seconds

Please give feedback!